OpenStudy (neverendingcycle):

What is the slope of the line passing through the points (0, 4) and (–8, –1)?

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (alexandervonhumboldt2):

\[\frac{ 4-(-1) }{ 0-(-8) }=?\]

OpenStudy (neverendingcycle):

5/-8

OpenStudy (neverendingcycle):

?

OpenStudy (alexandervonhumboldt2):

no